OpenStudy (anonymous):

this seems like a complicated one. can someone help me with this question? If the pH of a 1.00-in. rainfall over 1400miles2 is 3.60, how many kilograms of sulfuric acid, H2SO4, are present, assuming that it is the only acid contributing to the pH? For sulfuric acid, Ka1 is very large and Ka2 is 0.012.

OpenStudy (abmon98):

H2SO4(l) --> 2H+(aq) + SO4^2-(aq) pH = -log10[H+] [H+] = 10-pH [H+] = 10-3.60 [H+] = 0.00025 = 2.5 × 10-4M There are 2 moles present i think we should multiply by 2. 1 inch is 0.0254metres change the 1400 miles into m^2 multiply the 0.0254 and 1400 when you change the unit of it. Number of moles=concentration*volume After you obtain the number of moles use the equation Number of moles=Mass/Molar Mass Dont forget to change the Molar mass unit into kg Mass=Molar Mass*Number of moles Mass:98/1000*number of moles

OpenStudy (abmon98):

(Change the unit of miles square to m^2) 3625983354.42m^2 is equivalent to 1400 mile^2 m^2*m^1=m^3 3625983354.42*0.0254=92099977.2023m^3 Change to dm^3 9.20999772 × 10^10*2.5*10^-4=23024994.3 Moles 23024994.3*98/1000=mass mass=2256449 Kg

OpenStudy (anonymous):

I'm glad you explained in detail of the changing of the unit step. I was about to get lost then found your second post :) thank you so much!!! it makes more sense now!